# Diffusion Posterior Sampling is Computationally Intractable

### v1

@@ C.2 Small noise level – Score of vertex distribution close to full score in vertex orthant
Let $d^{\prime}=O(d)$. Consider $g_{s}$ and $g$ as in Definition 3.1. We have that for $x\in\{\pm 1\}^{d}$ such that $s$ is closest to $x_{1,\dots,d}$ among points in {±$\{\pm 1\}^{d}$, for the $\sigma$-smoothed versions $h_{s}=g_{s}*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$ of $g_{s}$ and $h=g*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$ of $g$, for $\frac{R^{2}}{1+\sigma^{2}}>C\log d$ for sufficiently large constant $C$,
@@
Let $d^{\prime}=O(d)$ and consider $g_{s}$ and $g$ as in Definition 3.1, and $x\in\mathbb{R}^{d+d^{\prime}}$. We have that for $s\in\{\pm 1\}^{d}$ such that $s$ is closest to $x_{1,\dots,d}$ among points in {±$\{\pm 1\}^{d}$, for $\sigma\geq\tau$, $\tau=\frac{1}{d^{C}}$ and $\varepsilon>\frac{1}{\operatorname{poly}(d)}$, for the $\sigma$-smoothed versions $h_{s}=g_{s}*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$ of $g_{s}$ and $h=g*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$ of $g$, for $\frac{R^{2}}{1+\sigma^{2}}>C\log d$ for sufficiently large constant $C$,
@@
Let $h_{s}$ be the $\sigma$-smoothed version of $g_{s}$, given by $h_{s}=g_{s}*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$. Let $\widetilde{s}\in\mathbb{R}^{d+d^{\prime}}$ be such that the first $d$ coordinates are given by $s$, and the remaining $d^{\prime}$ coordinates are 00. We have
@@ C.3 ReLU Network approximation of $\sigma$-smoothed Scores of Product Distributions
Since $f$ is piecewise linear, we can define $f$ as follows: there exists $-\infty=\gamma_{0}<\gamma_{1}<\gamma_{2}<\dots<\gamma_{D-1}=\gamma_{D}=+\infty$ such that
@@
This proves that $g(x)=f(x)$ for $x\in\mathbb{R}$ and we only need to design neural network to represent $g$. By employing one neuron for $a_{1}x+b_{1}$ and $D-1$ neurons for $\text{ReLU}((a_{i}-a_{i-1})(x-\gamma_{i-1}))$, and aggregating their outputs, we obtain the function $g$. There are $O(D)$ parameters in total, and each parameter is bounded by $O(\beta)$ in absolute value. ∎
@@ C.4 ReLU network for Score at Small smoothing level
For any $10<\alpha<$ 1, there exists a ReLU network $h:\mathbb{R}^{d}\to\mathbb{R}^{d}$ with $O(d/\alpha)$ parameters, constant depth, and weights bounded by $O(1/\alpha)$ such that
@@
[…]
@@
[…]
@@
[…]
@@
$h(x)_{i}=\begin{cases}\widetilde{h}_{i}(x_{i}-r(x)_{i}\cdot R)&\text{for $i\leq d$}\\ \text{switch}(\widetilde{h}_{i,1}(x_{i}),f(r(x))_{i-d})+\text{switch}(\widetilde{h}_{i,-1}(x_{i}),-f(r(x))_{i-d})&\text{for $i>d$}\\ \end{cases}$
@@
For $r(x)\in\{\pm 1\}^{d}$, consider the score $s_{\sigma,r(x)}$ of $g_{\sigma,r(x)}$, the $\sigma$-smoothed version of the distribution $g_{r(x)}$ centered at $\widetilde{r(x)}\in\mathbb{R}^{d+d^{\prime}}$, as described in Definition 3.1, where $\widetilde{r(x)}$ has the first $d$ coordinates given by $r(x)$, and the remaining coordinates set to 00.
@@
$\operatorname{\mathbb{P}}_{x\sim g_{\sigma}}\left[r(x)\not\in\{\pm 1\}^{d}\right]\lesssim de^{-\frac{(R-\alpha)^{2}}{2}}\lesssim e^{-\frac{R^{2}}{4}}$
@@
$\operatorname*{\mathbb{E}}_{x\sim g_{\sigma}}\left[\|s_{\sigma}(x)\|^{2}\cdot 1_{r(x)\not\in\{\pm 1\}^{d}}\right]$ $\leq\sqrt{\operatorname*{\mathbb{E}}_{x\sim g_{\sigma}}\left[\|s_{\sigma}(x)\|^{4}\right]\cdot\operatorname{\mathbb{P}}_{x\sim g_{\sigma}}\left[r(x)\not\in\{\pm 1\}^{d}\right]}$
@@ C.6 Large Noise Level - Distribution is close to a mixture of Gaussians
[…]
@@
$f_{\rho}(y)=\frac{1}{2^{d}}\sum_{x\in\{\pm 1\}^{d}}w_{\sqrt{\rho^{2}+1}}(y-R\cdot\widetilde{x})$
where $\widetilde{x}\in\mathbb{R}^{d+d^{\prime}}$ has the first $d$ coordinates given by $x$, and the last $d^{\prime}$ coordinates 00. Then, for $\frac{\rho^{2}}{\varepsilon^{2}(1+\rho^{2})}>C\log d$ for sufficiently large constant $C$, we have
@@ Appendix E Upper Bound
$\operatorname{\mathbb{P}}_{y}\left[\operatorname{\mathbb{P}}_{w}\left[\left\|y-w\right\|\leq 10\gamma\sqrt{m+\log(1/\delta)}\ \middle|\ y\right]\geq(c\gamma)^{m}\cdot\delta^{m/2+1}\right]\geq 1-\delta.$
@@
$\operatorname{\mathbb{P}}_{w\sim q}\left[\|w\|_{2}^{2}>\frac{Cm}{\delta}\right]<\frac{\delta}{3}.$
@@
$\operatorname{\mathbb{P}}[\exists\theta\in S:w\in\theta]\geq 1-\frac{\delta}{3}.$
@@
$S^{\prime}:=\{\theta\in S\mid\operatorname{\mathbb{P}}_{w}[w\in\theta]>\frac{\delta}{3N}\}.$
@@
$\operatorname{\mathbb{P}}_{w}\left[\forall\theta\in S^{\prime}:w\notin\theta\right]\leq\operatorname{\mathbb{P}}[\forall\theta\in S:w\notin\theta]+\operatorname{\mathbb{P}}[{\bigvee_{\theta\in S\setminus S^{\prime}}w\in\theta}]\leq\frac{\delta}{3}+N\cdot\frac{\delta}{3N}\leq\frac{2\delta}{3}.$
@@
$\operatorname{\mathbb{P}}_{y}\left[y\in S^{+}\right]$ $=\operatorname{\mathbb{P}}_{w\sim q,z\sim\mathcal{N}(0,I_{m})}\left[w+\beta z\in S^{+}\right]$
$\geq\operatorname{\mathbb{P}}_{w\sim q,z\sim\mathcal{N}(0,I_{m})}\left[(\exists\theta\in S^{\prime}:w\in\theta)\wedge(\left\|z\right\|\leq 8\sqrt{m})\right]$
$\geq 1-\operatorname{\mathbb{P}}_{w}\left[\forall\theta\in S^{\prime}:w\notin\theta\right]-\operatorname{\mathbb{P}}_{z\sim\mathcal{N}(0,I_{m})}\left[\left\|z\right\|^{2}>64(m+\log\frac{1}{\delta})\right]$
@@
$\operatorname{\mathbb{P}}_{z\sim\mathcal{N}(0,I_{m})}\left[\left\|z\right\|^{2}>64(m+\log\frac{1}{\delta})\right]<\frac{\delta}{3}.$
@@
$\operatorname{\mathbb{P}}_{y}\left[y\in S^{+}\right]\geq 1-\delta.$
This implies that with $1-\delta$ probability over $y$, there exists a cell $\theta\in S$ such that $\left\|{y}-\tilde{t}\right\|\leq 10\beta$ and $\operatorname{\mathbb{P}}_{w}[w\in\theta]\geq\frac{\delta}{3N}\geq\delta\cdot\Theta(\sqrt{\delta}\beta)^{m}$.
@@
Let $\mathcal{P}$ be the distribution that couples true distribution $\mathcal{D}$ over $(x,y)$ and the output distribution of the posterior sampler $\widehat{p}_{|y}$. Rigorously, we define $\mathcal{P}$ over $(x,\widehat{x},y)\in X\times X\times Y$ with density $p^{\mathcal{P}}$ such that $p^{\mathcal{P}}(x,y)=p^{\mathcal{D}}(x,y)$, $p^{\mathcal{P}}(\widehat{x}\mid y)=\widehat{p}_{\mid y}(\widehat{x})$. Similarly, we let $\widetilde{\mathcal{P}}$ over $(x,\widehat{x},y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{\alpha}$ be the joint distribution between the unconditional sampler over $(x,\widehat{x})$ and the measurement process $\mathcal{D}$ over $(x,y)$. Then by the definition of unconditional samplers, we have
$\operatorname{\mathbb{P}}_{x,\hat{x}\sim\widetilde{\mathcal{P}}}\left[\left\|x-\widehat{x}\right\|\geq\tau\right]\leq\delta.$
Therefore, to prove the correctness of the algorithm, we only need to show that there exists a $\widehat{\mathcal{P}}$ over $(x,\widehat{x},y)$ such that $\widetilde{\mathcal{P}}(\widehat{x}\mid y)=\widehat{p}_{\mid y}(\widehat{x})$ and $\textup{{TV}}(\widehat{\mathcal{P}},\widetilde{\mathcal{P}})\leq\delta$. By Lemma H.9,
$\operatorname{\mathbb{P}}_{\widetilde{\mathcal{P}}^{\prime}}\left[\left\|Ax-y\right\|^{2}\geq 4\beta^{2}(m+\log\frac{1}{\delta})\right]\leq\frac{\delta}{4}.$
@@
$\left|\log r(\widehat{x})\right|$ $\leq\sup_{x}\frac{|{\left\|Ax-y\right\|^{2}-\left\|A\widehat{x}-y\right\|^{2}}|}{2\beta^{2}}$
@@
By our setting of $\tau$, we have $81-\delta/8<r(\widehat{x})<1+\delta/$ 8.
@@
$\widehat{p}_{\mid y}(\widehat{x})$ $=\frac{r(\widehat{x})\cdot\int p^{\widetilde{\mathcal{P}}^{\prime}}(x,\widehat{x})e^{\frac{-\|A{x}-y\|^{2}}{2\beta^{2}}}\operatorname{d}\!{x}}{(1\pm\frac{\delta}{8})\int p^{\widetilde{\mathcal{P}}^{\prime}}(x,\widehat{x})e^{\frac{-\|A{x}-y\|^{2}}{2\beta^{2}}}\operatorname{d}\!{x}\operatorname{d}\!{\widehat{x}}}$
@@
$\operatorname{\mathbb{P}}_{x,\widehat{x}\sim\widehat{\mathcal{P}}}\left[\left\|x-\widehat{x}\right\|\geq\tau\right]\leq\frac{3\delta}{4}+\delta\leq\frac{7\delta}{4}.$
@@
$\geq\frac{1}{2}\operatorname{\mathbb{P}}_{x\sim\mathcal{X}}\left[\left\|Ax-y\right\|\leq 10\sqrt{m+\log(1/\delta)}\beta\right]\cdot e^{-\frac{100(m+\log(1/\delta))\beta^{2}}{2\beta^{2}}}$
$=\frac{1}{2}\operatorname{\mathbb{P}}_{x\sim\mathcal{X}}\left[\left\|Ax-y\right\|\leq 10\sqrt{m+\log(1/\delta)}\beta\right]\cdot\delta e^{-50m}$
By Lemma H.11, $\operatorname*{\mathbb{E}}_{x\sim\mathcal{X}}[{\left\|Ax\right\|_{2}^{2}}]=O(m)$. By Lemma E.1, we have that for $81-\delta/$ 8 probability over $y$, for some $c>0$,
$\operatorname{\mathbb{P}}_{x\sim\mathcal{X}}\left[\left\|Ax-y\right\|\leq 10\sqrt{m+\log(1/\delta)}\beta\right]\geq(c\beta)^{m}\cdot\delta^{m/2+1}.$
@@
$\operatorname{\mathbb{P}}_{y\sim\mathcal{Y}}\left[q(y)\geq(c\beta)^{m}\cdot\delta^{m/2+2}\right]\geq 1-\frac{\delta}{8}.$
@@
[Rejection sampling terminates in $\operatorname{\mathbb{P}}\left[\text{Rejection sampling terminates in }\frac{\log(1/\delta)}{\delta^{2}}\left(\frac{C}{\beta\sqrt{\delta}}\right)^{m}\text{rounds}\right]\geq 1-\frac{\delta}{4}.$

### v2

@@ C.2 Small noise level – Score of vertex distribution close to full score in vertex orthant
Let $d^{\prime}=O(d)$. Consider $g_{s}$ and $g$ as in Definition 3.2. We have that for $x\in\{\pm 1\}^{d}$ such that $s$ is closest to $x_{1,\dots,d}$ among points in $\{\pm 1\}^{d}$, for the $\sigma$-smoothed versions $h_{s}=g_{s}*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$ of $g_{s}$ and $h=g*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$ of $g$, for $\frac{R^{2}}{1+\sigma^{2}}>C\log d$ for sufficiently large constant $C$,
@@
Let $d^{\prime}=O(d)$ and consider $g_{s}$ and $g$ as in Definition 3.2, and $x\in\mathbb{R}^{d+d^{\prime}}$. We have that for $s\in\{\pm 1\}^{d}$ such that $s$ is closest to $x_{1,\dots,d}$ among points in $\{\pm 1\}^{d}$, for $\sigma\geq\tau$, $\tau=\frac{1}{d^{C}}$ and $\varepsilon>\frac{1}{\operatorname{poly}(d)}$, for the $\sigma$-smoothed versions $h_{s}=g_{s}*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$ of $g_{s}$ and $h=g*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$ of $g$, for $\frac{R^{2}}{1+\sigma^{2}}>C\log d$ for sufficiently large constant $C$,
@@
Let $h_{s}$ be the $\sigma$-smoothed version of $g_{s}$, given by $h_{s}=g_{s}*\mathcal{N}(0,\sigma^{2}I_{d+d^{\prime}})$. Let $\widetilde{s}\in\mathbb{R}^{d+d^{\prime}}$ be such that the first $d$ coordinates are given by $s$, and the remaining $d^{\prime}$ coordinates are 0. We have
@@ C.3 ReLU Network approximation of $\sigma$-smoothed Scores of Product Distributions
Since $f$ is piecewise linear, we can define $f$ as follows: there exists $\infty=\gamma_{0}<\gamma_{1}<\gamma_{2}<\dots<\gamma_{D-1}=\gamma_{D}=+\infty$ such that
@@
This proves that $g(x)=f(x)$ for $x\in\mathbb{R}$ and we only need to design neural network to represent $g$. By employing one neuron for $a_{1}x+b_{1}$ and D1D-1 neurons for $\text{ReLU}((a_{i}-a_{i-1})(x-\gamma_{i-1}))$, and aggregating their outputs, we obtain the function $g$. There are $O(D)$ parameters in total, and each parameter is bounded by $O(\beta)$ in absolute value. ∎
@@ C.4 ReLU network for Score at Small smoothing level
For any $0<\alpha<1$, there exists a ReLU network $h:\mathbb{R}^{d}\to\mathbb{R}^{d}$ with $O(d/\alpha)$ parameters, constant depth, and weights bounded by $O(1/\alpha)$ such that
@@
This is a piecewise linear function, where the derivative of each piece is bounded by $\frac{1}{\alpha}$, the value of the transition points are at most $\alpha$ in absolute value, and $|h|$ itself is bounded by $1$. Thus, by Lemma C.9, we can represent the function $h(x)=\left(g(x_{1}),\dots,g(x_{d})\right)$ using $O(d/\alpha)$ parameters, with each parameter’s absolute value bounded by $O(1/\alpha)$. Moreover, clearly $h(x)_{i}=\frac{x_{i}}{|x_{i}|}$ for all $i\in[d]$ whenever $|x_{i}|\geq\frac{1}{C}$. ∎
@@
Consider any function $\text{switch}:\mathbb{R}^{d+1}\to\mathbb{R}^{d}$ such that for $x\in\mathbb{R}^{d}$, $y\in\mathbb{R}$, with $|x_{i}|\leq T$ for all $i\in[d]$,
$\text{switch}(x,y)=\begin{cases}x&\text{if $y=1$}\\ 0&\text{if $y=-1$}\end{cases}$
@@
For $i\in d+[d^{\prime}]$, we will let $\widetilde{h}_{i,1}$ be the ReLU network that implements the approximation to the score $\widetilde{s}_{\sigma,i,-1}$ of $\widetilde{g}_{\sigma,i,-1}=(\frac{w_{1}\cdot\text{comb}_{\varepsilon}}{\int w_{1}(x)\cdot\text{comb}_{\varepsilon}(x)dx})*\mathcal{N}(0,\sigma^{2})$, and we will let $\widetilde{h}_{i,-1}$ implement the approximation to the score of $\left(\frac{w_{1}(x)\cdot\text{comb}_{\varepsilon}\left(x-\varepsilon/2\right)}{\int w_{1}(x)\cdot\text{comb}_{\varepsilon}(x-\varepsilon/2)dx}\right)*\mathcal{N}(0,\sigma^{2})$, as given by Lemma C.11. By the Lemma, for every $i\in d+[d^{\prime}]$ and $j\in\{\pm 1\}$, we have
@@
[…]
@@
For $r(x)\in\{\pm 1\}^{d}$, consider the score $s_{\sigma,r(x)}$ of $g_{\sigma,r(x)}$, the $\sigma$-smoothed version of the distribution $g_{r(x)}$ centered at $\widetilde{r(x)}\in\mathbb{R}^{d+d^{\prime}}$, as described in Definition 3.2, where $\widetilde{r(x)}$ has the first $d$ coordinates given by $r(x)$, and the remaining coordinates set to 0.
@@
$\operatorname*{\mathbf{Pr}}_{x\sim g_{\sigma}}\left[r(x)\not\in\{\pm 1\}^{d}\right]\lesssim de^{-\frac{(R-\alpha)^{2}}{2}}\lesssim e^{-\frac{R^{2}}{4}}$
@@
$\operatorname*{\mathbb{E}}_{x\sim g_{\sigma}}\left[\|s_{\sigma}(x)\|^{2}\cdot 1_{r(x)\not\in\{\pm 1\}^{d}}\right]$ $\leq\sqrt{\operatorname*{\mathbb{E}}_{x\sim g_{\sigma}}\left[\|s_{\sigma}(x)\|^{4}\right]\cdot\operatorname*{\mathbf{Pr}}_{x\sim g_{\sigma}}\left[r(x)\not\in\{\pm 1\}^{d}\right]}$
@@ C.6 Large Noise Level - Distribution is close to a mixture of Gaussians
Let $C$ be a sufficiently large constant. Let $g^{j}(x)\propto\prod_{i=1}^{d}w_{1}(x_{i})\cdot\prod_{i=d+1}^{d^{\prime}}w_{1}(x_{i})\cdot\text{comb}_{\varepsilon}(x_{i}-\phi_{i,j})$ be the pdf of a distribution on $\mathbb{R}^{d+d^{\prime}}$ with shifts $\phi_{i,j}$. Consider a mixture of discrete $d$-dimensional Gaussians, given by the pdf
@@
[…]
where $\widetilde{x}\in\mathbb{R}^{d+d^{\prime}}$ has the first $d$ coordinates given by $x$, and the last $d^{\prime}$ coordinates 0. Then, for $\frac{\rho^{2}}{\varepsilon^{2}(1+\rho^{2})}>C\log d$ for sufficiently large constant $C$, we have
@@ Appendix E Upper Bound
$\operatorname*{\mathbf{Pr}}_{y}\left[\operatorname*{\mathbf{Pr}}_{w}\left[\left\|y-w\right\|\leq 10\gamma\sqrt{m+\log(1/\delta)}\ \middle|\ y\right]\geq(c\gamma)^{m}\cdot\delta^{m/2+1}\right]\geq 1-\delta.$
@@
$\operatorname*{\mathbf{Pr}}_{w\sim q}\left[\|w\|_{2}^{2}>\frac{Cm}{\delta}\right]<\frac{\delta}{3}.$
@@
$\operatorname*{\mathbf{Pr}}[\exists\theta\in S:w\in\theta]\geq 1-\frac{\delta}{3}.$
@@
$S^{\prime}:=\{\theta\in S\mid\operatorname*{\mathbf{Pr}}_{w}[w\in\theta]>\frac{\delta}{3N}\}.$
@@
$\operatorname*{\mathbf{Pr}}_{w}\left[\forall\theta\in S^{\prime}:w\notin\theta\right]\leq\operatorname*{\mathbf{Pr}}[\forall\theta\in S:w\notin\theta]+\operatorname*{\mathbf{Pr}}[{\bigvee_{\theta\in S\setminus S^{\prime}}w\in\theta}]\leq\frac{\delta}{3}+N\cdot\frac{\delta}{3N}\leq\frac{2\delta}{3}.$
@@
$\operatorname*{\mathbf{Pr}}_{y}\left[y\in S^{+}\right]$ $=\operatorname*{\mathbf{Pr}}_{w\sim q,z\sim\mathcal{N}(0,I_{m})}\left[w+\beta z\in S^{+}\right]$
$\geq\operatorname*{\mathbf{Pr}}_{w\sim q,z\sim\mathcal{N}(0,I_{m})}\left[(\exists\theta\in S^{\prime}:w\in\theta)\wedge(\left\|z\right\|\leq 8\sqrt{m})\right]$
$\geq 1-\operatorname*{\mathbf{Pr}}_{w}\left[\forall\theta\in S^{\prime}:w\notin\theta\right]-\operatorname*{\mathbf{Pr}}_{z\sim\mathcal{N}(0,I_{m})}\left[\left\|z\right\|^{2}>64(m+\log\frac{1}{\delta})\right]$
@@
$\operatorname*{\mathbf{Pr}}_{z\sim\mathcal{N}(0,I_{m})}\left[\left\|z\right\|^{2}>64(m+\log\frac{1}{\delta})\right]<\frac{\delta}{3}.$
@@
$\operatorname*{\mathbf{Pr}}_{y}\left[y\in S^{+}\right]\geq 1-\delta.$
This implies that with $1-\delta$ probability over $y$, there exists a cell $\theta\in S$ such that $\left\|{y}-\tilde{t}\right\|\leq 10\beta$ and $\operatorname*{\mathbf{Pr}}_{w}[w\in\theta]\geq\frac{\delta}{3N}\geq\delta\cdot\Theta(\sqrt{\delta}\beta)^{m}$.
@@
[…]
$\operatorname*{\mathbf{Pr}}_{x,\hat{x}\sim\widetilde{\mathcal{P}}}\left[\left\|x-\widehat{x}\right\|\geq\tau\right]\leq\delta.$
[…]
$\operatorname*{\mathbf{Pr}}_{\widetilde{\mathcal{P}}^{\prime}}\left[\left\|Ax-y\right\|^{2}\geq 4\beta^{2}(m+\log\frac{1}{\delta})\right]\leq\frac{\delta}{4}.$
@@
[…]
@@
By our setting of $\tau$, we have $1-\delta/8<r(\widehat{x})<1+\delta/8$.
@@
[…]
@@
$\operatorname*{\mathbf{Pr}}_{x,\widehat{x}\sim\widehat{\mathcal{P}}}\left[\left\|x-\widehat{x}\right\|\geq\tau\right]\leq\frac{3\delta}{4}+\delta\leq\frac{7\delta}{4}.$
@@
$\geq\frac{1}{2}\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{X}}\left[\left\|Ax-y\right\|\leq 10\sqrt{m+\log(1/\delta)}\beta\right]\cdot e^{-\frac{100(m+\log(1/\delta))\beta^{2}}{2\beta^{2}}}$
$=\frac{1}{2}\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{X}}\left[\left\|Ax-y\right\|\leq 10\sqrt{m+\log(1/\delta)}\beta\right]\cdot\delta e^{-50m}$
By Lemma H.11, $\operatorname*{\mathbb{E}}_{x\sim\mathcal{X}}[{\left\|Ax\right\|_{2}^{2}}]=O(m)$. By Lemma E.1, we have that for $1-\delta/8$ probability over $y$, for some $c>0$,
$\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{X}}\left[\left\|Ax-y\right\|\leq 10\sqrt{m+\log(1/\delta)}\beta\right]\geq(c\beta)^{m}\cdot\delta^{m/2+1}.$
@@
$\operatorname*{\mathbf{Pr}}_{y\sim\mathcal{Y}}\left[q(y)\geq(c\beta)^{m}\cdot\delta^{m/2+2}\right]\geq 1-\frac{\delta}{8}.$
@@
$\operatorname*{\mathbf{Pr}}\left[\text{Rejection sampling terminates in }\frac{\log(1/\delta)}{\delta^{2}}\left(\frac{C}{\beta\sqrt{\delta}}\right)^{m}\text{ rounds}\right]\geq 1-\frac{\delta}{4}.$